OpenStudy (anonymous):

Find the x and y intercept of 4x+6y= -18

OpenStudy (anonymous):

First, to make this question easier, put the equation into the form y=mx+b by subtracting x then dividing by 6

OpenStudy (anonymous):

y-intecrepts: x=0 x-intercepts: y=0

OpenStudy (anonymous):

that is plug 0 in for x and 0 in for y

OpenStudy (anonymous):

Alternate forms: y = -2x/3 -3 4x + 6y + 18 = 0 2(2x + 3y) = -18 integer solution: x = 3n, y = -2n -3 solution for the variable y: y = 1/3 (-2 x - 9)

OpenStudy (triciaal):

x-intercept when y = 0 y -intercept when x = 0